Probability

5. The discrete random variable Z has the following probability distribution 2 0.2 4 P(Z) 0.1 0.25 0.05 0.3 Which of the following is FALSE? A) P(Z < 2) 0.55 B) P(Z 2 4)- 0.45 C) P(Z 4)=0.9 D) P(Z3)-0.05 6. A random sample of 100 first-year students was selected to determine the average GPA they achieved at a university. A 95% confidence interval for the average GPA of the first-year students was 5.2 < H < 6.8 based on sample statistics. The confidence can be interpreted as A) if all possible samples of 100 are drawn from the population, 95% of the confidence intervals developed will show that the sample mean will fall between 5.2 and 6.8 if all possible samples of 100 are taken from the population, 95% of the confidence intervals developed will show that the population mean is between 5.2 and 6.8 if all possible samples of 100 are taken from the population, 95% of the intervals developed will include the population mean 95% of first-year students have GPA between 5.2 and 6.8 B) C) D) 7. Which of the following is TRUE? A) The closer the correlation coefficient is to zero, the stronger the linear B) relationship between the two variables The correlation coefficient must lie in the range -1 to 1 The range of possible values for the coefficient of determination is -1 to C) D) Both B) and C) are correct.
Answer #1

Ans.5 - (B)

P(z>= 4) = P(Z=4)+P(Z=5) =0.3+0.1=0.4

Ans.6- (C)

X% Confidence interval states that if all samples of 100 are taken from population,X% of the intervals developed will contain true population parameter.

Ans.7- (B)

Correlation coefficient always lies between -1 and 1 .All the other options are incorrect.